# Reliable and efficient inverse analysis using physics-informed neural networks with distance functions and adaptive weight tuning

### v1

@@ 2.4 Positivity enforcement for inverse analysis
$\kappa\simeq\tilde{\kappa}\coloneqq h_{+}(\hat{\kappa}).$ (33)
@@ 3 Numerical results
This section presents the results of several numerical experiments. Throughout the experiments, we employed MLPs with a depth of 5 and a width of 64, initialized by the Glorot method [77]. This choice is based on prior studies demonstrating its effectiveness across a broad range of problems [39, 70, 63, 78]. The Adam optimizer [73] was used for training, with exponential decay rates of 0.9 and 0.999 for the first and second moment estimates, respectively.
As is well known, the choice of the activation function significantly affects the convergence and stability. Common choices include the hyperbolic tangent function and the rectified linear unit (ReLU, $\max(0,\cdot)$). However, the ReLU and some of its variants [79, 80, 81] are generally unsuitable for PINNs because their higher-order derivatives vanish over a wide range of inputs. This limitation hinders the accurate computation of higher-order derivatives of the solution, leading to inaccurate PDE residual calculations [82]. Several studies have also reported numerical results that the ReLU leads to poor convergence in PINNs [83, 84]. Recent research has proposed the rectified power unit ($\text{RePU}_{p}=\max(0,\cdot)^{p},p\in\mathbb{Z}_{\geq 2}$) as an alternative [46, 12]. However, RePU networks are prone to vanishing gradients for inputs less than 1 and exploding gradients for inputs greater than 1, particularly in deeper networks. Indeed, Sukumar and Srivastava [46] and Samaniego et al. [12] utilized RePU networks with a depth of 3, which is relatively shallow in the literature. Our preliminary experiments also confirmed that the RePU network with increased depth (4, 5, and 6) exhibited vanishing and exploding gradient issues, leading to instability. Consequently, we evaluated and compared the hyperbolic tangent, Sigmoid Linear Unit (SiLU) [85, 86], and Gaussian Error Linear Unit (GELU) [87] activation functions. These functions are $C^{\infty}$-continuous and have non-vanishing higher-order derivatives, which are essential for properly calculating the PDE residuals.
Furthermore, several techniques have been proposed to enhance PINN performance, including adaptive activation functions [83, 88, 89], gradient enhancement [25], and self-adaptive soft attention mechanisms [90, 91]. However, based on the authors’ experience, the accuracy improvements achieved by these methods are often marginal and highly sensitive to hyperparameter selection (see Appendix A for numerical results). Hence, we focus on the fundamental aspects of PINN, such as boundary condition treatment and adaptive weight tuning.
We have implemented all the experiments using TensorFlow [92] 111 https://github.com/ShotaDeguchi/PINN_HardBC_DynNorm. .
@@ 3.1 Forward problem: Poisson equation with mixed boundary conditions
(a) Soft imposition (weight parameter: $\lambda_{\text{BC}}=1$).
(b) Hard imposition (normalization order: $m=1$).
@@
$g_{D}$ =$=\begin{cases}\sin\left(\pi x\right)&\text{on}\quad\Gamma_{3},\\ 0&\text{on}\quad\Gamma_{2}\cup\Gamma_{4},\end{cases}$ (35)
@@
Figure 6 presents reference solution, approximate solution, absolute error, and the first derivative of the approximate solution along the negative $y$-direction, so that one can observe how well the Dirichlet and Neumann boundary conditions are satisfied. While a visual comparison of the second column suggests that soft and hard impositions yield similar solutions, the third and fourth columns reveal a significant difference. The soft imposition fails to properly enforce the prescribed Dirichlet and Neumann boundary conditions, whereas the hard imposition does so effectively, as evidenced by the absolute error and derivative plots. The relative $L^{2}$ error further emphasizes the importance of boundary condition treatment. The error of the hard imposition is $3.13\times 10^{-3}$, which is an order of magnitude lower than that of the soft imposition, $2.88\times 10^{-2}$. Figure 7 shows the relative $L^{2}$ error of the approximate solution with different activation functions and boundary condition impositions. We observe from Figure 7 (a) that the soft imposition exhibits a plateau regardless of the activation function, which is known as pitfall of soft boundary condition enforcement [39]. This behavior is likely due to the presence of multiple competing objectives, which makes it difficult for the network to learn the correct solutions [63]. In contrast, the hard imposition demonstrates superior convergence, achieving a relative error of $\mathcal{O}(10^{-3})$ (Figure 7 (b)). Comparing the convergence behavior of different activation functions, we find that both soft and hard impositions converge more rapidly with SiLU and GELU than with tanh. This should be attributed to the non-saturating property of SiLU and GELU, which allows the network to learn more effectively. Moreover, GELU exhibits slightly faster convergence than SiLU, which is likely due to its steeper gradient around the origin. Furthermore, Table 1 summarizes the relative error with different activation functions and boundary condition impositions. One can observe that the hard imposition consistently outperforms the soft imposition across all activation functions, confirming the effectiveness of the hard boundary condition imposition in PINN.
@@
Figure 8 presents reference solution, approximate solution, absolute error, and the first derivative of the approximate solution along the negative $y$-direction to visually realize how well the Neumann condition is satisfied. Again, the soft imposition fails to meet the prescribed boundary conditions, whereas the hard imposition effectively handles the inhomogeneous Neumann condition. We also observed that the soft imposition suffers from the same ’plateau’ issue as in the homogeneous Neumann condition case, while the hard imposition is free from this trouble (figures not shown). Table 2 summarizes the relative error with different activation functions and boundary condition impositions. These results further confirm the effectiveness and applicability of the hard BC imposition. Based on these results, GELU demonstrates superior performance among the three activation functions considered. Therefore, we will employ GELU as the activation function in the subsequent experiments.
@@ 3.2 Inverse problem: shear-driven cavity flow of an incompressible fluid
where $\bm{u}=(u,v)^{\top}$ is the velocity, $p$ is the pressure, and $\mathrm{Re}$ is the Reynolds number, set to 1,000 and 5,000 in this study. Boundary conditions (39) and (40) correspond to the driving and the stationary walls, respectively (definitions of $\Gamma_{1}$$\Gamma_{4}$ are consistent with those in Equations (18)–(21)). The reference solution was obtained using the finite difference method on a staggered Arakawa B-type grid [93]. Advection was discretized with a third-order upwind scheme [94, 95], while diffusion and pressure gradient were approximated using second-order central differences. Velocity and pressure were coupled via Chorin’s projection method [96]. The spatial resolution was set to $\Delta x=\Delta y=5\times 10^{-3}$, which was confirmed to be sufficiently fine for the Reynolds numbers considered. The time step was selected as $\Delta t=1\times 10^{-3}$ in accordance with the stability criterion for explicit time marching [97, 98]. The steady state was assumed when the following criterion was met:
$\frac{\|u^{(n+1)}-u^{(n)}\|}{\|u^{(n)}\|}\leq\varepsilon,$ (41)
@@
where $\varepsilon$ is the tolerance, set to $1\times 10^{-8}$. We have confirmed that the obtained numerical solution is in good agreement with the previous studies [99, 100] and is treated as the reference solution for the inverse analysis. With the obtained reference solutions, we formulate the inverse problem of identifying the Reynolds number $\mathrm{Re}$ and the pressure $p$ from limited observations of the velocity field $\bm{u}$ (no pressure data is provided to the network). Collocation points were randomly drawn from the domain with $N_{\text{PDE}}=4,096$ and $N_{\text{Data}}=256$, with an additional $N_{\Gamma}=256\times 4$ boundary points for the soft imposition. The learning rate was initially set to $10^{-3}$, and an exponential decay schedule was applied with a decay factor of 0.9 every 2,000 iterations. Similar results were observed made with other learning rate schedules, including cosine decay. However, the exponential decay schedule demonstrated superior performance in terms of convergence and stability compared to cosine decay and constant learning rate schedules. For both soft and hard boundary condition impositions, bias-corrected dynamic normalization was employed to adaptively adjust the weights for data loss (Equation (8)) and divergence loss (the residual of Equation (37)). The decay rates were chosen based on our previous studies [47, 74], which indicated that higher decay rates are more effective. While higher decay rates might typically lead to slower convergence, the introduced bias correction (Equation (32)) effectively mitigates this issue, which is a unique feature of our method to differentiate from similar approaches, such as [30, 24, 70, 31]. For the soft imposition approach, the weight for boundary loss was also adaptively adjusted by dynamic normalization. We also evaluated tanh, SiLU, and GELU activation functions, observing similar trends to those presented in Section 3.1. Consequently, we only present the results obtained using the GELU activation function.
(a) Reference solution, reconstructed velocity, and inferred pressure.
(b) Identification of the Reynolds number (mean and standard deviation over 5 i.i.d. runs).
[…]
(a) Reference solution, reconstructed velocity, and inferred pressure.
(b) Identification of the Reynolds number (mean and standard deviation over 5 i.i.d. runs).
[…]
@@
Figure 9 and 10 present the reconstructed velocity, inferred pressure, and estimated Reynolds number for $\mathrm{Re}=1,000$ and $5,000$ cases, respectively. The importance of boundary condition treatment is highlighted, especially by comparing the velocity norm. Specifically, we observe that the soft imposition fails to recover the velocity field, specifically due to the violation of the prescribed no-slip boundary condition on $\Gamma_{1}\cup\Gamma_{2}\cup\Gamma_{4}$. Consequently, the soft imposition approach also fails to estimate the pressure distribution; the output is inaccurate and physically unrealistic, and the Reynolds number. Given that PINN performance in inverse analysis is highly dependent on boundary condition treatment [46], this failure, primarily attributed to improper boundary condition enforcement, underscores the critical impact of accurate approximate solution construction on inverse analysis.
@@
However, the sole application of the hard imposition is still insufficient. In the third column of Figure 9 and 10, it can be observed that the hard imposition approach precisely satisfies the boundary conditions (Equations (39)–(40)); however, the reconstructed velocity field resembles that of a highly viscous fluid, which is not consistent with the references at the corresponding Reynolds numbers. This discrepancy arises from the inability of the hard imposition alone to balance data and PDE losses, thereby preventing the effective integration of observed data into the network.
In contrast, the combination of hard imposition and dynamic normalization achieves a balance between multiple losses, leading to accurate reconstruction of the velocity field and pressure distribution, as demonstrated in the fourth column of Figure 9 and 10. The Reynolds number is also accurately estimated for both cases. These results convincingly demonstrate the robustness and effectiveness of the hard boundary condition imposition and adaptive weight tuning for inverse problems using PINNs.
@@ 3.3 Inverse problem: incompressible flow around an obstacle
Method Estimated kinematic viscosity [$\mathrm{m^{2}/s}$]
@@
Following [9], we consider an unsteady flow over an obstacle. The governing equations read as:
@@
where $\bm{u}=(u,v)^{\top}$ is the velocity [$\mathrm{m/s}$], $p$ is the pressure [$\mathrm{Pa}$], $\rho$ is the density [$\mathrm{kg/m^{3}}$], and $\nu$ is the kinematic viscosity [$\mathrm{m^{2}/s}$]. We set the density and the kinematic viscosity to $\rho=1\ [\mathrm{kg/m^{3}}]$ and $\nu=2\times 10^{-3}\ [\mathrm{m^{2}/s}]$. The domain is $4\ [\mathrm{m}]\times 1\ [\mathrm{m}]$ rectangle with a square obstacle of side length $0.2\ [\mathrm{m}]$ located at $(1,0.5)$. No-slip conditions are applied on the obstacle, top, and bottom boundaries. A parabolic velocity profile with a maximum velocity of $1\ [\mathrm{m/s}]$ is given on the inlet boundary, and a zero-gradient condition is imposed on the outlet. The corresponding Reynolds number is $\mathrm{Re}\sim 100$. The reference solution was obtained through a FDM simulation, employing the same computational scheme as in Section 3.2, with a spatial resolution of $\Delta x=\Delta y=1\times 10^{-2}\ [\mathrm{m}]$ and a time step of $\Delta t=1\times 10^{-3}\ [\mathrm{s}]$. The simulation was executed until the flow reached a periodic state, and a 5-second segment of this periodic flow was extracted as the reference solution. Similar to the cavity flow problem, we consider an inverse problem of identifying the kinematic viscosity $\nu$ and the pressure $p$ from limited observations of the velocity field $\bm{u}$. This study differs from the previous works [9, 70] by considering the entire computational domain, rather than focusing solely on the wake region. This highlights the effect of different boundary condition imposition techniques on the performance of the inverse analysis using PINN. Moreover, in this study, the distance is built based on R-functions [50, 51, 52], which makes it possible to detect interior edges (cylinder boundary), as well as the outer boundary. This treatment was infeasible with the approaches described in [2, 39, 45, 34, 66], as they primarily focused on detecting outer boundaries defined by convex hulls. Collocation points were randomly drawn with $N_{\text{PDE}}=1,024$ and $N_{\text{Data}}=64$ every $0.1\ [\mathrm{s}]$, resulting in a total of $N_{\Omega\times\mathcal{I}}=52,224$ and $N_{\text{Data}\times\mathcal{I}}=3,264$. For soft imposition, $N_{\Gamma_{\text{inlet}}\times\mathcal{I}}=N_{\Gamma_{\text{outlet}}\times\mathcal{I}}=5,151$, $N_{\Gamma_{\text{top}}\times\mathcal{I}}=N_{\Gamma_{\text{bottom}}\times\mathcal{I}}=20,451$, $N_{\Gamma_{\text{cylinder}}\times\mathcal{I}}=4,284$ points were added for the boundary condition enforcement. For this problem, we only employed GELU, as it demonstrated the best performance in previous sections. The Adam optimizer with a learning rate of $1\times 10^{-3}$ was used.
@@
Figure 11 (a) shows the reference, reconstructed velocity, and inferred pressure fields, as well as the estimated kinematic viscosity for the square obstacle case. While the soft imposition demonstrates reasonable agreement with the reference in the wake region, it faces significant challenges in accurately capturing the flow pattern near the obstacle. In particular, the velocity field exhibits non-physical flow separation upstream of the object, with its magnitude being underestimated both near the obstacle and within the wake region. Furthermore, the pressure field shows a non-physical distribution, with the peak point shifted upstream from the obstacle. Conversely, the combination of hard imposition and dynamic normalization captures the flow structure accurately, including the region around the cylinder. Although a slight underestimation of the velocity magnitude is observed in the wake, the discrepancy is considerably smaller than that of the soft imposition. The pressure field, which was not provided to the network, is also inferred with high accuracy, with the peak pressure point correctly located on the obstacle surface. It is important to note that, due to the nature of the incompressible flow, the pressure is determined only up to an additive constant [101]. Therefore, the slight shift observed in the inferred pressure under hard imposition and dynamic normalization is not a concern.
@@
The results of the kinematic viscosity estimation are shown in Figure 11 (b) and summarized in Table 3. Figure 11 (b) reveals that the soft imposition yields highly unstable viscosity estimates, rendering it unsuitable for practical applications. As can be seen in Table 3, the soft imposition, even with dynamic normalization, fails to accurately identify the kinematic viscosity, undershooting the true value by orders of magnitude. In contrast, the combination of hard imposition and dynamic normalization provides stable and accurate viscosity estimates, closely matching the true value. Notably, the estimation error in the viscosity is only 1.9% for $\beta=0.99$ and 0.4% for $\beta=0.999$, where $\beta$ is the decay rate of dynamic normalization. Furthermore, training did not converge in the absence of dynamic normalization, as indicated in Table 3. These findings underscore the critical role of hard boundary condition imposition and dynamic normalization in enhancing the performance of PINNs in inverse analysis. In addition, Figure 12 illustrates the evolution of the adaptive weights assigned to the data and divergence losses. One can observe that the weights are adaptively adjusted from the early stages of training, highlighting the effectiveness of bias-corrected dynamic normalization in balancing the data and divergence losses.
@@
(a) Distribution of collocation points.
(b) Reconstructed velocity field.
@@
Method Estimated kinematic viscosity [$\mathrm{m^{2}/s}$]
@@
The methodology is further extended to the flow around a heart-shaped obstacle. The governing equations are the same as in the previous case (Equations (42) and (43)). The domain is a $4\ [\mathrm{m}]\times 1\ [\mathrm{m}]$ rectangle, and the square cylinder is replaced with a heart-shaped obstacle. The same kinematic viscosity and boundary conditions are applied, and the geometric parameters are chosen so that the same Reynolds number is maintained. The reference solution was obtained using a P2/P1 finite element method [102] on a triangular mesh. The P2/P1 interpolation was chosen to satisfy the inf-sup condition, ensuring the stability of the numerical solutions [103]. The mesh was progressively refined towards the obstacle, resulting in minimum, maximum, and average element lengths of $3.5\times 10^{-3}\ [\mathrm{m}]$, $2.4\times 10^{-2}\ [\mathrm{m}]$, $1.2\times 10^{-2}\ [\mathrm{m}]$, respectively. The problem was solved with varying mesh resolutions, confirming the convergence of the numerical results with the current resolution. The flow exhibits periodic, yet asymmetric behavior, attributed to the obstacle’s geometry. The obtained numerical solution is used as the reference for the inverse analysis. The same number of collocation points and identical settings for the learning rate, dynamic normalization, and activation function were used as in the square obstacle case.

### v3

@@ 2.2.3 Positivity enforcement for inverse analysis
[…]
@@ 3 Numerical results
This section presents the results of several numerical experiments. Throughout the experiments, we employed MLPs with a depth of 5 and a width of 64, initialized by the Glorot method [Glorot2010]. This choice is based on prior studies demonstrating its effectiveness across a broad range of problems [Sun2020Surrogate, Jin2021NSFnets, Rohrhofer2023IEEE]. The Adam optimizer [Kingma2015Adam] was used for training, with exponential decay rates of 0.9 and 0.999 for the first and second moment estimates, respectively.
As is well known, the choice of the activation function significantly affects the convergence and stability. Common choices include the hyperbolic tangent function and the rectified linear unit (ReLU, $\max(0,\cdot)$). However, the ReLU and some of its variants [Maas2013LeakyReLU, He2015, Clevert2016ELU] are generally unsuitable for PINNs because their higher-order derivatives vanish over a wide range of inputs. This limitation hinders the accurate computation of higher-order derivatives of the solution, leading to inaccurate PDE residual calculations [Deguchi2021Doboku]. Several studies have also reported numerical results that the ReLU leads to poor convergence in PINNs [Jagtap2020GAAF, Haghighat2021SolidMechanics]. Recent research has proposed the rectified power unit ($\text{RePU}_{p}=\max(0,\cdot)^{p},p\in\mathbb{Z}_{\geq 2}$) as an alternative [Sukumar2022MixedBC, Samaniego2020Energy]. However, RePU networks are prone to vanishing gradients for inputs less than 1 and exploding gradients for inputs greater than 1, particularly in deeper networks. Indeed, Sukumar and Srivastava [Sukumar2022MixedBC] and Samaniego et al. [Samaniego2020Energy] utilized RePU networks with a depth of 3, which is relatively shallow in the literature. Our preliminary experiments also confirmed that the RePU network with increased depth (4, 5, and 6) exhibited vanishing and exploding gradient issues, leading to instability. Consequently, we evaluated and compared the hyperbolic tangent, Sigmoid Linear Unit (SiLU) [Elfwing2017SiLU, Ramachandran2017Swish], and Gaussian Error Linear Unit (GELU) [Hendrycks2016GELU] activation functions. These functions are $C^{\infty}$-continuous and have non-vanishing higher-order derivatives, which are essential for properly calculating the PDE residuals.
Furthermore, several techniques have been proposed to enhance PINN performance, including adaptive activation functions [Jagtap2020GAAF, Jagtap2020LAAF], gradient enhancement [Yu2022GradEnhc], and self-adaptive soft attention mechanisms [McClenny2023SA]. However, based on the authors’ experience, the accuracy improvements achieved by these methods are often marginal and highly sensitive to hyperparameter selection (see Appendix A for numerical results). Hence, we focus on the fundamental aspects of PINN, such as boundary condition treatment and adaptive weight tuning.
We have implemented all the experiments using TensorFlow [Abadi2015TensorFlow]  [^1] https://github.com/ShotaDeguchi/PINN_HardBC_DynNorm. .
@@ 3.1 Forward problem: Poisson equation with mixed boundary conditions
(a) Soft imposition (weight parameter: $\lambda_{\text{BC}}=1$)
(b) Hard imposition (normalization order: $m=1$)
@@
$g_{D}$ $=\begin{cases}\sin\left(\pi x\right)&\text{on}\quad\Gamma_{3},\\ 0&\text{on}\quad\Gamma_{2}\cup\Gamma_{4},\end{cases}$ (36)
@@
Figure 6 presents reference solution, approximate solution, absolute error, and the first derivative of the approximate solution along the negative $y$-direction, so that one can observe how well the Dirichlet and Neumann boundary conditions are satisfied. Although the second column may appear similar, the third and fourth columns reveal a significant difference. The soft imposition fails to properly enforce the prescribed Dirichlet and Neumann boundary conditions, whereas the hard imposition does so effectively, as evidenced by the absolute error and derivative plots. The relative $L^{2}$ error further emphasizes the importance of boundary condition treatment. The error of the hard imposition is $3.13\times 10^{-3}$, which is an order of magnitude lower than that of the soft imposition, $2.88\times 10^{-2}$. Figure 7 shows the relative $L^{2}$ error of the approximate solution with different activation functions and boundary condition impositions. We observe from Figure 7 (a) that the soft imposition exhibits a plateau regardless of the activation function, which is known as pitfall of soft boundary condition enforcement [Sun2020Surrogate]. This behavior is likely due to the presence of multiple competing objectives, making it difficult for the network to learn the correct solutions [Rohrhofer2023IEEE]. In contrast, the hard imposition demonstrates superior convergence, achieving a relative error of $\mathcal{O}(10^{-3})$ (Figure 7 (b)). Comparing the convergence behavior of different activation functions, we find that both soft and hard impositions converge more rapidly with SiLU and GELU than with tanh. This should be attributed to the non-saturating property of SiLU and GELU, which allows the network to learn more effectively. Moreover, GELU exhibits slightly faster convergence than SiLU, which is likely due to its steeper gradient around the origin. Furthermore, Table 2 summarizes the relative error with different activation functions and boundary condition impositions. One can observe that the hard imposition consistently outperforms the soft imposition across all activation functions, confirming the effectiveness of the hard boundary condition imposition in PINN.
@@
(a) Soft imposition (weight parameter: $\lambda_{\text{BC}}=1$)
(b) Hard imposition (normalization order: $m=1$)
Figure 9: Poisson equation with mixed boundary conditions: inhomogeneous Neumann condition case. Relative $L^{2}$ error of the approximate solution with different activation functions. The solid line represents the mean, and the shaded area represents the standard error over 5 independent runs with i.i.d. initialization. Soft imposition quickly plateaus regardless of the activation function.
@@
Figure 8 presents reference, approximate solution, absolute error, and the first derivative of the approximate solution along the negative $y$-direction to visually realize how well the Neumann condition is satisfied. Again, the soft imposition fails to meet the prescribed boundary conditions, whereas the hard imposition effectively handles the inhomogeneous Neumann condition. One can also observe that the soft imposition suffers from the same ’plateau’ issue as in the homogeneous Neumann condition case, while the hard imposition is free from this trouble (Figure 9). Table 3 summarizes the relative error with different activation functions and boundary condition impositions. These results further confirm the effectiveness and applicability of the hard BC imposition. Based on these results, GELU demonstrates superior performance among the three activation functions considered. Therefore, we will employ GELU in the subsequent experiments.
@@ 3.2 Inverse problem: shear-driven cavity flow of an incompressible fluid
where $\bm{u}=(u,v)^{\top}$ is the velocity, $p$ is the pressure, and $\mathrm{Re}$ is the Reynolds number, set to 1,000 and 5,000 in this study. Boundary conditions (40) and (41) correspond to the driving and the stationary walls, respectively (definitions of $\Gamma_{1}$$\Gamma_{4}$ are consistent with those in Equations (19)–(22)). The reference solution was obtained using the finite difference method on a staggered Arakawa B-type grid [Arakawa1977Grid]. Advection was discretized with a third-order upwind scheme [Kawamura1984, Kawamura1986], while diffusion and pressure gradient were approximated using second-order central differences. Velocity and pressure were coupled via Chorin’s projection method [Chorin1968Projection]. The spatial resolution was set to $\Delta x=\Delta y=5\times 10^{-3}$, which was confirmed to be sufficiently fine for the Reynolds numbers considered. The time step was selected as $\Delta t=1\times 10^{-3}$ in accordance with the stability criterion for explicit time marching [Courant1967, Neumann1950]. The steady state was assumed when $\|u^{(n+1)}-u^{(n)}\|/\|u^{(n)}\|\leq 1\times 10^{-8}$ was met. We have confirmed that the obtained numerical solution is in good agreement with the previous studies [Ghia1982Cavity, Erturk2005Cavity] and is treated as the reference solution for the inverse analysis. With the obtained reference solutions, we formulate the inverse problem of identifying the Reynolds number $\mathrm{Re}$ and the pressure $p$ from limited observations of the velocity field $\bm{u}$ (no pressure data is provided to the network). Collocation points were randomly drawn from the domain with $N_{\text{PDE}}=4,096$ and $N_{\text{Data}}=256$, with an additional $N_{\Gamma}=256\times 4$ boundary points for the soft imposition. The learning rate was initially set to $10^{-3}$, and an exponential decay schedule was applied with a decay factor of 0.9 every 2,000 iterations. For both soft and hard boundary condition impositions, bias-corrected dynamic normalization was employed to adaptively adjust the weights for data loss (Equation (8)) and divergence loss (the residual of Equation (38)). The decay rates were chosen based on our previous studies [Deguchi2023DynNorm, Deguchi2023Keisan], which indicated that higher decay rates are more effective. While higher decay rates might typically lead to slower convergence, the introduced bias correction (Equation (33)) effectively mitigates this issue, which is a unique feature of our method to differentiate from similar approaches, such as [Wang2021GradPath, Maddu2022InvDirichlet, Jin2021NSFnets, Wang2023ExpertGuide]. For the soft imposition approach, the weight for boundary loss was also adaptively adjusted by dynamic normalization.
(a) Reference solution, reconstructed velocity, and inferred pressure
(b) Relative $L^{2}$ error of velocity
(c) Identification of Reynolds number
Figure 10: Shear-driven cavity flow of an incompressible fluid: $\mathit{Re=1,000}$ case. The application of hard imposition alone remains ineffective. Accurate solutions are achieved by combining hard imposition and dynamic normalization. In panels (b) and (c), the solid (or dotted) line represents the mean over 5 i.i.d. runs, and the shaded area represents the standard error in (b) and the standard deviation in (c).
(a) Reference solution, reconstructed velocity, and inferred pressure
(b) Relative $L^{2}$ error of velocity
(c) Identification of Reynolds number
Figure 11: Shear-driven cavity flow of an incompressible fluid: $\mathit{Re=5,000}$ case. Similar to the $\mathrm{Re}=1,000$ case, the sole application of hard imposition fails. The combination of hard imposition and dynamic normalization achieves accurate inverse analysis. As in Figure 10, panels (b) and (c) depict the mean (solid or dotted line) with shaded regions representing the standard error in (b) and the standard deviation in (c) over 5 i.i.d. runs.
@@
Figures 10 and 11 present the reconstructed velocity, inferred pressure, estimated Reynolds number, and relative $L^{2}$ error of velocity field for the $\mathrm{Re}=1,000$ and $5,000$ cases, respectively. The importance of boundary condition treatment is highlighted, especially by comparing the velocity norm. Specifically, we observe that the soft imposition fails to recover the velocity field, particularly due to the violation of the prescribed no-slip boundary condition on $\Gamma_{1}\cup\Gamma_{2}\cup\Gamma_{4}$. Consequently, the soft imposition approach also fails to estimate the pressure distribution; the output is inaccurate and physically unrealistic. Given that PINN performance in inverse analysis is highly dependent on boundary condition treatment [Sukumar2022MixedBC], this failure, primarily attributed to improper boundary condition enforcement, underscores the critical impact of accurate approximate solution construction on inverse analysis.
@@
However, we stress that the sole application of the normalized distance field remains insufficient. In the third column of Figures 10 (a) and 11 (a), it can be observed that the hard imposition approach precisely satisfies the boundary conditions (Equations (40)–(41)); however, the reconstructed velocity field resembles that of a highly viscous fluid, which is not consistent with the references at the corresponding Reynolds numbers. This discrepancy arises from the inability of the hard imposition alone to adequately balance data and PDE losses, thereby preventing the effective integration of observed data into the network. Notably, Figures 10 (b) and 11 (b) further indicate that the plateau issue persists in both the soft and hard approaches within this inverse analysis.
In contrast, the combination of hard imposition and dynamic normalization achieves a balance between multiple losses, leading to accurate reconstruction of the velocity field and pressure distribution, as demonstrated in the fourth column of Figures 10 (a) and 11 (a). The relative error in the velocity field is significantly reduced, not only when compared to soft imposition but also in comparison to the results obtained with hard imposition alone (Figures 10 (b) and 11 (b)). The Reynolds numbers are also accurately estimated for both cases, as depicted in Figures 10 (c) and 11 (c). Notably, the successful inverse analysis achieved by our proposed framework at a Reynolds number of $5,000$, without the need for curriculum training, represents a significant advancement in the application of PINNs to challenging regimes. It is well-known that directly training PINNs for high Reynolds number flows often leads to instability and convergence to erroneous solutions, necessitating the use of staged curriculum learning [Krishnapriyan2021FailureModes]. Our results, demonstrating accurate flow field inference without curriculum training, underscores the robustness and efficacy achieved by the integration of hard boundary condition enforcement and bias-corrected adaptive weight tuning.
@@ 3.3 Inverse problem: incompressible flow around an obstacle
Method Estimated kinematic viscosity $\mathrm{m^{2}/s}$]
@@
Following [Raissi2019PINN], we consider an unsteady flow over an obstacle. The governing equations read as:
@@
[…]
@@
Figure 12 (a) shows the reference, reconstructed velocity, and inferred pressure fields, as well as the estimated kinematic viscosity for the square obstacle case. While the soft imposition demonstrates reasonable agreement with the reference in the wake region, it faces significant challenges in accurately capturing the flow pattern near the obstacle. In particular, the velocity field exhibits non-physical flow separation upstream of the object, with its magnitude being underestimated both near the obstacle and within the wake region. Furthermore, the pressure field shows a non-physical distribution, with the peak point shifted upstream from the obstacle. Conversely, the combination of hard imposition and dynamic normalization captures the flow structure accurately, including the region around the cylinder. Although a slight underestimation of the velocity magnitude is observed in the wake, the discrepancy is considerably smaller than that of the soft imposition. The pressure field, which was not provided to the network, is also inferred with high accuracy, with the peak pressure point correctly located on the obstacle surface. It is important to note that, due to the nature of the incompressible flow, the pressure is determined only up to an additive constant [ChorinMarsden1993]. Therefore, the slight shift observed in the inferred pressure under hard imposition and dynamic normalization is not a concern.
@@
The results of the kinematic viscosity estimation are shown in Figure 12 (b) and summarized in Table 4. Figure 12 (b) reveals that the soft imposition yields highly unstable viscosity estimates, rendering it unsuitable for practical applications. As can be seen in Table 4, the soft imposition, even with dynamic normalization, fails to accurately identify the kinematic viscosity, undershooting the true value by orders of magnitude. In contrast, the integration of hard imposition and dynamic normalization provides stable and accurate viscosity estimates, closely matching the true value. Notably, the estimation error in the viscosity is only 1.9% for $\beta=0.99$ and 0.4% for $\beta=0.999$, where $\beta$ is the decay rate of dynamic normalization. Furthermore, training did not converge in the absence of dynamic normalization, as indicated in Table 4. These findings underscore the critical role of hard boundary condition imposition and dynamic normalization in enhancing the performance of PINNs in inverse analysis. In addition, Figure 13 illustrates the evolution of the adaptive weights assigned to the data and divergence losses. One can observe that the weights are adaptively adjusted from the early stages of training, highlighting the effectiveness of bias-corrected dynamic normalization in balancing the data and divergence losses.
@@
(a) Distribution of collocation points
(b) Reconstructed velocity field
@@
Method Estimated kinematic viscosity $\mathrm{m^{2}/s}$]
@@
[…]